OpenStudy (anonymous):

What is the value of the expression? 100 – 23 + 8 · 4 A. 124 B. 126 C. 400 D. 408

OpenStudy (anonymous):

I'm 99% sure it's A. Just checking.

OpenStudy (anonymous):

@Hero @Whitemonsterbunny17 @SolomonZelman @midhun.madhu1987 @cwrw238

OpenStudy (gahm8684):

it gives me 109

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

^

OpenStudy (anonymous):

is it a ?

OpenStudy (cwrw238):

do the last step 8 times 4 first then work from left to right 100 - 23 + 32 = 77 + 32 = 109

OpenStudy (anonymous):

so whats the answer

OpenStudy (cwrw238):

if that is 8 times 4 then 109 is correct times is usually written * though

OpenStudy (cwrw238):

why do you think its a?

OpenStudy (cwrw238):

you sure you've got the question right?

OpenStudy (anonymous):

yah

OpenStudy (cwrw238):

all i can say as i've said earlier then its 109 if thats a multiplication at the end multiply is done before add/subtract you heard of PEDMAS?

OpenStudy (anonymous):

YAH the answer was a lol

OpenStudy (anonymous):

There's no way the answer was A.

OpenStudy (anonymous):

it was lol

OpenStudy (anonymous):

You had to of listed the wrong answers or the wrong question.